Determine the length of curve 𝑦 = log(sec 𝑥) between 0 ≤ 𝑥 ≤

𝜋

4


1
Expert's answer
2021-05-07T09:06:14-0400

Given f(x) = y = log(s"\because" ec(x)) on [0,"\frac{\pi}{4}"]


The arc length is given by L = "\int_{a}^{b} \sqrt{(f'(x))^2 +1 } \,dx"

we have f(x) = log(sec(x))

"\implies" f'(x) = "\frac{1}{sec(x)} \cdot" (sec(x) "\cdot" tan(x)) = tan(x)


"\therefore" L = "\int_{0}^{\frac{\pi}{4}} \sqrt{(tan(x))^2 +1 }dx"


= "\int_{0}^{\frac{\pi}{4}} \sqrt{tan^2(x) +1 }dx"


= "\int_{0}^{\frac{\pi}{4}} \sqrt{sec^2(x) }dx" ( "\because" sec2(x) - tan2(x) = 1)

= "\int_{0}^{\frac{\pi}{4}} sec(x) dx"

= [ log( sec(x) +tan(x) )]"|^\frac{\pi}{4}_0"


= [log(tan("\frac{\pi}{4}") + sec("\frac{\pi}{4}")] - [log(tan(0) + sec(0)]


= [log(1+ "\sqrt{2}" )] - [log(1)]


"\therefore" L = log(1+ "\sqrt{2}" )